[Hallc_running] [New Logentry] RC Daily Summary August 25, 2025
ole at jlab.org
ole at jlab.org
Mon Aug 25 09:45:01 EDT 2025
Logentry Text:
--
Hall C data taking over the last 24 hours went very smoothly as far as beam delivery was concerned. We logged 18.22 ABU hours (76% efficiency). However, we had some glitches on our end, resulting in about 4 hours of probably unusable data. After the SHMS polarity switch during DAY, two quads were inadvertently left at the wrong polarity (https://logbooks.jlab.org/entry/4449971). Fortunately, Steve caught this right away at the start of SWING and got us back on track (https://logbooks.jlab.org/entry/4449993). Runs 25223–25226 are affected. Then, early during OWL, the presale GUI was accidentally closed, and the wrong GUI was started afterwards, resulting in wrong prescale factors for runs 25245 and 25246. This was recognized and corrected quickly.
Reminders/lessons learned:
- After a polarity change, double check all magnet settings (AND the replay later!). For all the magnets of a given spectrometer, the "POL" column in the summary GUI must indicate the same thing, either "E" or "P", never a mix of "E" and "P". The alarm handler setpoints appear to be set correctly after a polarity change, so any magnet-related alarm should be cause for concern.
- The correct prescale GUI can be started with "go_prescale_SHMS" from the hccoda account.
We have finished the z=0.90 pi+ theta_pq=2.0 setting (pages 43-45 in the run plan) and, as of about 6:30am, are taking data for the z=0.67 pi+ theta_pq=2.0 kinematics (pages 46-49 in the run plan, ETA 2-3 shifts). We tentatively plan to do the BCM calibration tomorrow.
Updates from Accelerator:
- Hall A finished the SBS run this morning around 7am and will be down for the rest of this run period (and beyond).
- No beam for some time during DAY because of the Hall B pass change.
- The troubleshooting of the Arc10 power supply for Hall D will happen over the next day or two and will be (mostly) not invasive to beam delivery to the other halls.
- There will be a maintenance day on Wednesday. Hall C will go into restricted access from about 9am-noon for a tour.
Shift summaries:
8/24 DAY: https://logbooks.jlab.org/entry/4449735
8/24 SWING: https://logbooks.jlab.org/entry/4449962
8/25 OWL: https://logbooks.jlab.org/entry/4450164
We will have the regular RC meeting at 2:30 pm today in the counting house conference room.
Zoom link: https://jlab-org.zoomgov.com/j/1618227115?pwd=8JkhFg6YNesVc3HNgWjwUSIKIu5SvP.1
Meeting ID: 161 822 7115
Passcode: 253463
Dial-in: 833 568 8864
---
This is a plain text email for clients that cannot display HTML. The full logentry can be found online at https://logbooks.jlab.org/entry/4450429
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20250825/ecc532f2/attachment.htm>
More information about the Hallc_running
mailing list